Four pencils and five pens cost $2.00.three pencils and four pens cost $1.58.find the cost of a pencil an the cost of a pen.
Answer by jorel1380(3719)

4m+5n=2.00
3m+4n=1.58
12m+15n=6.00
12m+16n=6.32
n=.32
m=.10
Pencils cost 10 cents; pens cost 32 cents..